OpenStudy (anonymous):

Simplify cotxtanx? Would the answer be 1?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Yes. Cotangent is equal to 1/tanx. So: \[\frac{ 1 }{ \tan x } \times \tan x = \frac{ \tan x }{ \tan x } = 1\]